DRAFT

\underline{1}

Table 1: Shielded pickup witness bunch data sets recorded from March, 2010 through November, 2011. Note that the various cloud-mitigating coatings (TiN, a-carbon, and diamond-like carbon) were swapped into the different radiation environments at 15E and 15W. Data sets with differing total radiation dose were obtained as well. Two vacuum chambers with a-carbon coatings were studied, providing information on production reproducibility.

Date  Species Beam Energy Bunch Current 15E/W Mitigation Bunch Spacing
(GeV)  (mA)  Technique  (ns)
03/27/2010 Positrons  5.3  5  W  a-carbon (1)  14-84
E  TiN
Electrons 5 W a-carbon (1) 14-70
E  TiN
05/09/2010 Positrons 2.1 3 W Al 4-140
E  a-carbon (2)
Electrons 3 W Al 4-20
E  a-carbon (2)
05/17/2010 Positrons 5.3 3 W Al 4-100
E  a-carbon (2)
Electrons 3 W Al 4-100
E  a-carbon (2)
05/19/2010 Electrons 2.1 1 W Al 4-120
E  a-carbon (2)
09/21/2010 Positrons 5.3 1,2,4,6,8,10 W TiN 14
E  a-carbon (2)
09/24/2010 Positrons 2.1 2,4,6 W TiN 14
E  a-carbon (2)
Electrons W TiN
E  a-carbon (2)
12/10/2010 Electrons 2.1 1,2,3,4,5,6,8,10 W TiN 14-84
E  a-carbon (2)
12/20/2010 Positrons 2.1 1,2,3,4,5,6,8,10 W TiN 56,84
E  a-carbon (2)
12/24/2010 Positrons 5.3 3,5 W TiN 14-84
E  a-carbon (2)
Electrons 3,5 W TiN 14-84
E  a-carbon (2)
04/07/2011 Positrons 5.3 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
Electrons 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
04/16/2011 Positrons 2.1 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
04/17/2011 Electrons 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
06/11/2011 Positrons 2.1 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
06/12/2011 Electrons 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
06/18/2011 Positrons 5.3 1,2,3,4,5,6,8,10 W TiN 14-98
E  DL carbon
Electrons 1,2,3,4,5,6,8,10 W TiN 14-84
E  DL carbon
06/27/2011 Positrons 4.0 1,2,3,4,5,6,8,10 W TiN 14-98
E  DL carbon
Electrons 2.1 1,2,3,4,5,6 W TiN 84
E  DL carbon
09/27/2011 Positrons 5.3 1,2,3,4,5,6,8 W a-carbon (2) 14-84
E  DL carbon
09/30/2011 Positrons 5.3 1,2,3,4,5,6,8 W a-carbon (2) 14-84
E  DL carbon
10/04/2011 Positrons 5.3 1,2,3,4,5,6,8 W a-carbon (2) 14-84
E  DL carbon
10/11/2011 Positrons 5.3 1,2,3,4,5,6,8 W a-carbon (2) 14-84
E  DL carbon
10/25/2011 Positrons 5.3 1,2,3,4,5,6,8 W a-carbon (2) 14-84
E  DL carbon
11/27/2011 Positrons 5.3 1,2,3,4,5,6,8,10 W a-carbon (2) 14-98
E  DL carbon
Electrons 1,2,3,4,5,6,8,10 W a-carbon (2) 14-84
E  DL carbon